[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jul 27, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jul 27 11:58:18 UTC 2017


Here is today's summary of NHC computer operations:
--- PS

----------

TAFB:
1.) Problems with labels and Danger graphic. Problems appear to be with  
GEMPAK Program. Matt Sardi is aware and he will be looking into.

2.) EPAC GRIDS SPF would not load on GRUNT. SPF GUI stayed grayed out. Matt  
Sardi to look into it.

3.) Atlantic NWPS running but not coming into GFE. Chris Mello will look at  
this first thing today. Jeff L sent email with details.
HSU:
1.) JTWC is still having problems downloading our data and some of our  
guidance is not coming is as normal. TSB is aware of it and spent all day  
yesterday investigating. Problem is Opah, a server at NCEP. We could not  
get NCO to resolve it yesterday. This has been elevated to highest priority  
today and if no action is taken we will involve the Director.

2.) NRL Data problems continue. Mike and Monica have been following this up  
but it is out of our hands. NRL is aware of it and working on it.

3.) HSU would like TSB to make higher priority to get G16 SWIR loop  
working. Noted. Monica will have this. But the other data flow problems  
above and NetApp migration have take precedence until resolved.

4.) Jack reporting CAVE sessions freezing. When this happens, restart.  
Also, open a TT with NCF and notify Chris Mello.
Other items:

NHC RFCs:
1.) NetApp Migration continues with big switching taking place today. This  
might impact us on different on fronts. Your patience and cooperation will  
be most appreciated. Please let us know of anything that does not look  
right ;)


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170727/daa8e434/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list